OpenStudy (anonymous):

Solve for the variable. a - 16 = -10

OpenStudy (anonymous):

Isolate the variable a by adding 16 to both sides of the equation

OpenStudy (anonymous):

a=6

OpenStudy (anonymous):

how you do solve it though?

OpenStudy (anonymous):

a - 16 = -10 + 16 = +16 ______________ a = 6

OpenStudy (anonymous):

first, we isolate "A" by adding 16 to both sides... then work it out as usual...